Monday, 23 November 2015

Gear on the move down

Today, Nick Moyes moved all the various technical gear and gyzmos down a floor inside the cathedral tower. This involved turning the whole system off so we apologise if you were getting a black screen today!
It took two of us about six hours to make the move down. Unplugging everything was no simple task and each lead had to be labelled to ensure it went back in the right socket!
Fortunately the wiring Nick had employed back in 2007 when the web cams were set up was long enough to be fed through the ringing room floor and down into the room below.
Our new home is bigger than the previous one and well supplied with sockets too!
Clearly there may well be some teething problems (and the pictures seemed to be somewhat blurred at first) but we hope to have these sorted out when we next do a maintenance abseil....please bear with us.
Nick M figures out the wiring in
its new location
Finally, as if by magic, a picture appears on the monitor. There's some tidying up yet to be done but at least the system seems to be working.
The monitor in position with an old display board from upstairs
adjacent. This will be replaced before the summer.
A problem with Stream 4 seems to be caused by some external factor interfering with the signal as it travels from the tower to the Silk Mill Museum or, more likely, from the museum across to the council's IT.
One possible cause is a new hotel being built further down Full Street behind the cathedral. 
The new hotel which may possibly be getting in the way of the signal
with the nave roof in the foreground

Sorting this out is beyond our control but we have reported the problem - so fingers crossed.
We are aware of the ongoing problems with the camera that looks horizontally above the nest (towards Jurys Inn) and we plan to address this before the start of the new breeding season.

The project team
Ps. Stream 3 is off air until the spring.

No comments: